If Adenine is present 20% in DNA then concentration of pyrimidine in DNA would be:   (A) 80%   (B)20%    (C)50%    (D)70% 

Grade:12th pass

1 Answers

Apurva Sharma
5 years ago
DNA constitutes: adenine(A), guanine(G),cytosine(C) and thymine(T)
Adenine and Guanine have PURINE bases
Thymine and Cytosine have PYRIMIDINE bases
 
Now ATQ,
%A = 20
And we know that amount in DNA of A is equal to T & of G is equal to C
Thus, %T =20
 
now, %G = %C, 
thus, %(G+C) = 100- %(A+T) =60%
thus %G =%C =30%
Therefore. Concentration of Pyrimidine: %(T+C) = 50%
 
so, the correct answer is option C.
 
Hope it cleared all the doubts.
Then please do give a thumbsup!😊👍